OpenStudy (anonymous):

can someone help me with 1 question.

OpenStudy (jlg030597):

yes

OpenStudy (anonymous):

Call: r the speed of jet traffic from Albuquerque to Atlanta t1 travel time from Albu. to Atlanta. Speed from Atlanta to Albu.: (r + r/18) = 19*r/18 t2 travel time from Atlanta to Albu. We have: 1270 = r*t1 = (19*r/18) t2 -> 18r*t1 = 19r* t2 -> -> 18*t1 = 19t2. This is the simplified expression..